EX 5:- 1 The molecular weight of carbon dioxide, CO2



, is 44. In an experiment the value



γ for CO



2 was found to be 1.3. Assuming that CO



2



is a perfect gas, calculate the gas



constant, R, and the specific heats at constant pressure and constant volume, Cp, Cv



(0.189 kJ/kg.K; 0.63kJ/kg.K; 0.819kJ/kg.K).

1
Expert's answer
2022-04-21T17:44:23-0400

"с_v = { i*R \over 2\mu}= {6*8,31 \over 2 * 0,044}(J)={6*8,31 \over 2 * 0,044}:1000= 0,566({kJ \over kgK})"

"с_p = { (I+2)*R \over 2\mu}= {8*8,31 \over 2 * 0,044}(J)={8*8,31 \over 2 * 0,044}:1000=0,755({kJ \over kgK})"

R="c_p-c_v=0,755-0,566=0,189({kJ \over kgK})"